UT 20130328 Technical Observing - TCS Software Tests & Active Optics Calibration for MODS1

Observer: JHill, CBiddick, MDeLaPena, TSargent (Tucson) DMiller (home)
Telescope Operator: SAllanson (LBTO)
Additional AO Support: JChristou (LBTO)
Telescope Support: JLittle (LBTO)
Instrument Support: TShih (LBTO)

Summary

Sky was very cloudy, but the seeing was good.

We spent the first 4 hours on software testing with emphasis on: testing patches, understanding binocular offsets, looking for stale coordinates from MCSPU, and understanding problems with moving hotspots.

We had significant problems with GCS and AzCams. Had to cold boot azcamserver to get things running again after 4 warm restarts didn't do the trick.

Doug M took over from software at 06:15 and finished calibrating the signs of Z12-Z21 for MODS1.

Lost the last 1.5 hours to really heavy clouds - sometimes 8.4m aperture is not enough.

Preparations

Details

Open

02:15 Open with fairly heavy broken clouds

TCS Software Testing (with a couple patches to Build 2013A)

Authorized LUCI1@LFBG and IRTC@RFBG (with RIP'd shell)

Binocular Offset Tests

02:20 SYNC Acquire presets on pointing star

02:30 SYNC Active presets on-axis --- seeing is 0.5 arcsec, at least on the left side

02:50 RFBG rotator hangs ??

03:03 Parallactic ACQUIRE on right

increment right IE by -40 arcsec

03:06 +40 IE increment on right ---> left star jumps by 1 arcsec

03:08 offset 35 0 RADEC ABS on right

03:09 offset -35 0 RADEC ABS on right

03:10 offset 0 0 RADEC ABS on right

03:11 Set leaking to zero, and restart PCS

Resend presets

03:16 increment right IE by -40 arcsec

03:17 increment right IE by +40 arcsec

03:19 offset 35 0 RADEC ABS on right

03:20 offset 0 0 RADEC ABS on right

03:21 increment right CA by -40 arcsec

03:22 increment right CA by +40 arcsec - we don't see the opposite side jump when leaking is zero

03:23 Set leaking to 10.0 and restart PCS

03:28 increment right IE by -40 arcsec - opposite star moves by 2 arcsec, and guiding is oscillating

03:32 increment right IE by +40 arcsec, and again the star oscillates

03:45 change presets to off-axis stars- acquisition offset misses and then oscillates for 5 cycles

03:52 increment right IE by -40 arcsec

increment right IE by +40 arcsec

04:01 put leak rate back to 2.5 and restart PCS

04:13 increment right IE by -40 arcsec

04:15 increment right IE by +40 arcsec - this made larger jump than on-axis at 2.5

NSI Drift Guiding Tests @LFBG

04:41 NSI preset to same star with 100 arcsec/hour drift rate

04:48 preset again

tcslog says ROIs are updating, but image suggests not

blue circle is updating correctly

04:52 resend preset - not NSI mode

04:57 resend preset NSI

IIF hung trying to abort preset - apparently a known bug that Chris has fixed in the next build.

05:11 Pointing star ....

AzCam wierdness......GCS is hanging waiting for the image on BOTH sides. Restarted multiple times.

Now with RA rate = 50

Updating ROI seems to correctly have a threshold of 10 pixels, but the y-value appears to stop updating before the ROI reaches the edge of the chip.

05:35 right preset succeeds -- left fails

05:56 restarting IIF, PCS, GCSL, GCSR along with AzCamg1

06:00 Sky is 75% cloudy

06:03 preset to a new star - authorized only at LUCI1@LFBG - probe can't move

AzCam on AGW1 is still hung......

06:07 stop GCSL, restart AzCamG and AzCam server

06:08 stop and restart PCS for Michele

06:10 Software people depart

Active Optics Calibration

06:18 Reconfigure for MODS1

06:24 Shutdown azcamserver and agw1azcamg for cold boot, also stopping IIF, PCS, GCSL, GCSR

06:38 Turn on MODS1 agw unit

06:42 Steve succeeds with a readGuideCam image from AGW1.

06:44 Preset to LUCI1 with swing arms in MODS configuration - we get an acquisition image.

06:45 Authorize as MODS1@LFBG

Turn on multi_step_focus script

06:49 ACQUIRE preset doesn't show a star - clouds are pretty thick

06:50 problems setting SX M2

06:55 ACTIVE preset off-axis ACT0294 - lovely moonlit clouds for testing the background removal by sextractor

Guide cam exposure time (ms): 659
WFS cam exposure time (ms): 17958
AGW position: 0.4472, -77.6187 (native: -0.199, -77.575)

07:01 Seeing is 0.6 to 0.7 on the guider, and we appear to be looking though ~3 mags of cap cloud. Initially sending low-order Zernikes plus Z22.

07:09 Doug online - turning on all the Z12-Z22 Zernikes.

Took a couple MODS images, but they are moon-illuminated flat fields more than star images.

07:23 HO Zernikes were applied, but Doug isn't sure why - he hadn't save the file yet.

07:25 Clouds wipe out the guide star and telescope walks away (the star was really gone - should GCS have recognized that it was gone in this much moon? FWHM of the blob it picked was 2.2 arcsec.)

Calibrating the signs of Z12-21 for MODS1

07:30 Active preset to AO1000 in NE, but still problems with thick and patchy clouds

07:54 Star lost to clouds

07:55 Resend preset ......

08:00 Star lost to clouds again - Steve says there are no holes left.

09:10 Guide window on GCSGUI stops updating, and the star slowly drifts off the hotspot.

09:30 Stop GCSL and resend the preset.

09:4x Doug trying to adjust MODS1 hotspot, but GCS won't update the hotspot it is using. ????

09:55 Reconfigure to LUC1@LFBG

10:08 ACQUIRE preset to pointing star

10:20 We're losing 8-10 magnitudes as the clouds are getting thicker.

Close

10:22 Close for HEAVY clouds.

-- JohnHill - 28 Mar 2013

Photo by Steve Allanson: SunPillar_20130327_184253_261.jpg:
SunPillar_20130327_184253_261.jpg
Topic revision: r6 - 29 Mar 2013, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback